I keept getting the totals for the per share wrong, please help

Pecan Theatre Inc. owns and operates movie theaters throughout Florida and Georgia. Pecan Theatre has declared the following annual dividends over a six-year period: 20Y1, $80,000; 20Y2, $90,000; 20Y3, $150,000; 20Y4, $150,000; 20Y5, $160,000; and 20Y6, $180,000. During the entire period ended December 31 of each year, the outstanding stock of the company was composed of 250,000 shares of cumulative, preferred 2% stock, $20 par, and 500,000 shares of common stock, $15 par.

Required:

1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of 20Y1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0".

Preferred Dividends Common Dividends

Year
Total
Dividends

Total

Per Share

Total

Per Share
20Y1 $   80,000 $ $ $ $
20Y2 90,000
20Y3 150,000
20Y4 150,000
20Y5 160,000
20Y6 180,000
$ $
Answer #1

In case of cumulative preference shares ,any dividend not paid on such shares in any years is carried to future years and are paid in preference to current year preferred and common stock dividend .

preferred dividend per year =par value *dividend rate

            = (250000*20)*2%

            = 5000000*2%

            = 100000

Preferred dividend common stock dividend Preferred dividend in arrear
Year Dividend Total Per share Total per share
20Y1 80000 80000 80000/250000=$ .32 0 0 100000-80000=20000
20Y2 90000 90000 90000/250000=$ .36 0 0 100000-90000=10000 current +20000last year =30000
20Y3 150000 30000 arrear +100000 current year =130000 130000/250000 =$ .52 per share 150000-130000=20000 20000/500000 =$ .04 per share 0
20Y4 150000 100000 100000/250000 =$ .40 per share 150000-100000=50000 50000/500000=$ .1 per share 0
20Y5 160000 100000 .40 160000-100000=60000 60000/500000 =$ .12 per share 0
20Y6 180000 100000 .40 180000-100000=80000 80000/500000=$ .16 per share 0
Total 810000 600000 $ 2.4 per share 210000 .42 per share